# Learning Spatial-Semantic Features for Robust Video Object Segmentation

### v1

@@ Abstract
Tracking and segmenting multiple similar objects with complex or separate parts in long-term videos is inherently challenging due to the ambiguity of target parts and identity confusion caused by occlusion, background clutter, and long-term variations. In this paper, we propose a robust video object segmentation framework equipped with spatial-semantic features and discriminative object queries to address the above issues. Specifically, we construct a spatial-semantic network comprising a semantic embedding block and spatial dependencies modeling block to associate the pretrained ViT features [18, 34] with global semantic features and local spatial features, providing a comprehensive target representation. In addition, we develop a masked cross-attention module to generate object queries that focus on the most discriminative parts of target objects during query propagation, alleviating noise accumulation and ensuring effective long-term query propagation. The experimental results show that the proposed method set a new state-of-the-art performance on multiple datasets, including DAVIS2017 test (89.1%), YoutubeVOS 2019 (88.5%), MOSE (75.1%), LVOS test (73.0%), and LVOS val (75.1%), which demonstrate the effectiveness and generalization capacity of the proposed method. We will make all source code and trained models publicly available.
@@ 1 Introduction
Video Object Segmentation (VOS) aims to track and segment target objects that are specified in the initial frame with mask annotations in a video sequence [47, 36, 40, 50]. This practice holds significant promise across various applications, particularly as the prevalence of video content surges in domains like autonomous driving, augmented reality, and interactive video editing [35, 2]. The main challenges faced by VOS include drastic target appearance change, occlusion, and identity confusion caused by similar objects and background clutter, which becomes more difficult, particularly in long-term videos.
@@
VOS methods [33, 11] typically perform video object segmentation by comparing the test frame with past frames. Specifically, they first use an association model to generate the correlated features of the test sample and target templates pixel-wise and then predict target masks based on the correlated features. The pixel-wise correlated features [44] contribute to accurate mask predictions. To account for targets that appear differently over time, some methods [50, 10] utilize a memory module to store these changing appearances. In addition, several recent methods [42, 14] introduce object queries to help distinguish different target objects for alleviating identity confusion.
@@
Despite the good performance on short-term test videos with simple-structured objects, existing methods do not work well in the following challenging scenarios. First, when dealing with target objects with multiple complex or separate parts caused by occlusion, background clutter, and shape complexity, existing methods often generate incomplete prediction masks. This is because the pixel-wise correlation adopted by existing methods mainly focuses on detailed spatial information at the pixel level and neglects the semantic information at the instance level. Second, although object query improves the ID association accuracy, it does not work well in sequences with dramatic target appearance changes. Existing methods [42, 14] update target queries relying on the entire predicted sample, which can introduce noise and errors and further lead to tracking failures, where targets are missed or their identities are swapped.
@@
Figure 1: Performance comparison in terms of $\mathcal{J}\&\mathcal{F}$ on various datasets. The proposed method achieves favorable performance against the state-of-the-art methods, including AOT_L [50], XMem [10], and Cutie [14] on all the datasets. All methods are trained using the YoutubeVOS and DAVIS datasets.
@@
To address the aforementioned issues, we propose a VOS framework to learn spatial-semantic features and generate discriminative object queries for robust VOS. To be specific, we construct a Spatial-Semantic Block comprising a semantic embedding module and a spatial dependencies modeling module to efficiently leverage the semantic information and local details of the pre-trained ViTs for VOS without training all the parameters of the ViT backbone. In addition, we develop a discriminative query to focus more on the representative target parts for more reliable target representation and query update, which can alleviate noise accumulation in query propagation and ensure robustness in long-term videos. We evaluate the proposed method extensively on five benchmarks, including DAVIS 2017 [36], Youtube VOS2018 [47], Youtube VOS2019, LVOS [22], and MOSE [17]. The experimental results show that our approach set new state-of-the-art performance on all these datasets including DAVIS2017 test (89.1%), YoutubeVOS 2019 (88.5%), MOSE (75.1%), and LVOS test (73.0%). In addition, the results achieved by only using the DAVIS and Youtube datasets for training show that our method generalizes well to different datasets, as shown in Figure 1.
@@
  • We propose a spatial-semantic network block to incorporate semantic information with spatial information for Video Object Segmentation. We use the cls token from a pre-trained ViT backbone to extract global semantic information for interaction with current features. In addition, we design a spatial dependencies modeling module to leverage the features of the visual transformer better. The proposed spatial-semantic block enables robust VOS performance in handling the target objects with multiple complex or separate parts.
  • We develop a discriminative query mechanism to capture the representative region of the target for better target representation learning and updating. This achieves significant improvement in VOS datasets, especially in long-term VOS datasets.
  • We conduct extensive experiments on five diverse datasets to demonstrate the effectiveness of the proposed method and comprehensive ablation studies to evaluate the contribution of each proposed component. We set a new state-of-the-art performance on all these datasets.
@@ 2 Related Work
Video Object Segmentation. Video Object Segmentation (VOS) aims to identify and separate objects within a video sequence from the background and from each other based on one annotated target frame, which is also known as semi-supervised VOS. Early VOS methods [4, 46, 41, 23, 31] use test-time learning to adapt pre-trained segmentation models for separating the specified targets online. Test-time learning of a deep model is time-consuming and significantly reduces the run speed. To avoid online learning, propagation-based methods [19, 16, 39] model the temporal correlations between adjacent frames with offline shift attention learning to propagate target masks from the previous frame to the current. Despite the promising performance, these methods suffer from error accumulation caused by inaccurate predictions. To achieve fast and efficient inference, matching-based methods [7, 24, 3, 49] identify targets by comparing the target template to the test image and predict target masks based on matching features. For instance, OML [7] and VideoMatch [24] conduct pixel matching between the nearest frames for mask prediction. Some methods [33, 50, 51, 10] further improve the matching-based VOS framework by introducing target memory to enrich target templates and developing advanced matching strategies to generate more comprehensive correlated features. For example, SimVOS [44] and JointFormer [32] use the ViT blocks [18] to model features and patch correspondence for object modeling jointly. In addition, some recent methods [42, 14] introduce object queries to enrich target representation for better ID identification.
Despite the rich target features and sufficient interactions, existing methods lack modeling of object-level semantic information, so they do not work well on dealing with objects with complex structures and long-term deformation. Unlike existing methods, the proposed method learns features and associates templates with test frames at both pixel and object levels achieved through a novel object-aware backbone and a pixel-and-query level correlation component, leading to robust VOS performance on complex and long-term videos.
@@
Query Transformer Networks. DETR [5] utilizes queries to represent target regions for object detection tasks. In the segmentation domain, some methods [9, 6] introduce masked cross-attention to use queries to represent targets for segmentation tasks. Motivated by the above methods, ISVOS [42] utilizes additional queries generated by Mask2Former [9] to represent instance information for target modeling, enhancing target representation in matching-based VOS methods. To achieve end-to-end training, Cutie [14] learns object-level information and facilitates interaction between object and pixel information, improving segmentation accuracy and efficiency. However, the query-based VOS methods suffer from query drift due to ineffective propagation. These methods model the entire target sample in the query, incorporating redundant information and degrading saliencies of the target object. Instead, the proposed method generates discriminative queries for target representation and query propagation, which is more robust to variations and preserves discriminative information.
@@
Adapters for Dense Tasks. Vit-Adapter [8] first addresses the problem of utilizing pre-trained models to boost detection and segmentation performance and proposes a dense adapter to exchange information with ViT blocks. To enhance spatial hierarchical features, ViT-CoMer [45] introduces a multi-receptive field feature pyramid module to provide multi-scale spatial features. However, those two methods require full fine-tuning for the best performance, which is inefficient, especially in the video domain. Meanwhile, they focus on spatial information enhancement while lacking semantic feature interaction between ViT blocks and multi-scale features.
@@
To better suit VOS tasks, our method features several design distinctions compared to existing adapters. First, we fuse the feature from ViT to multi-scale features and do not need extra modules to transmit information to ViT blocks, which is more efficient. Second, the proposed fusion block associates semantic priors and local dense features. Third, the proposed method freezes the ViT blocks during training and does not need full fine-tuning.
@@ 3 Proposed Algorithm
Our method aims to learn comprehensive target features that contain semantic, spatial, and discriminative information for video object segmentation. This helps us cope with complex target appearance variations and ID confusion between target objects with similar appearances in long-term videos. To this end, we propose a spatial-semantic feature learning network that first embeds semantic information from a trained ViT network with multi-scale features from a trainable CNN network. Then, we model spatial dependencies upon the fused feature, enabling spatial-semantic features for VOS. In addition, to ensure that the target query retains discriminative information during long-term appearance changes, we develop a discriminative query propagation module to capture and update the local representations in the target object.
@@ 3.1 Overall Framework
Figure 2(a) shows the overall framework of the proposed method, consisting of two core modules: a spatial-semantic feature generation module and a pixel-and-query dual-level target association module. Given a test frame $\mathbf{X}_{t}\in\mathbb{R}^{3\times H\times W}$, the proposed method predicts the mask of all specified objects within it based on the initial frame $\mathbf{X}_{init}\in\mathbb{R}^{3\times H\times W}$ and its corresponding annotation mask $\mathbf{M}_{init}\in\mathbb{R}^{H\times W}$. First, the feature generation module takes the test frame $\mathbf{X}_{t}$ as input and generates spatial-semantic features. Then, the target association module associates the spatial-semantic features with the reference samples using pixel- and query-wise correlation to generate the correlated features. Finally, a decoder predicts the mask based on these correlated features.
To better represent video targets that vary over frames, we develop a dual-level memory module consisting of the base features generated by the feature generation module and target queries that capture more global information. Both types of memory are updated every few frames using the online predicted masks, enabling adaptive target representations for effective target association.
Figure 2: Architecture of the proposed method. (a) The overall framework of the proposed method. (b) Implementation of the spatial-semantic block. Semantic priors and location details are fused by semantic embedding and spatial dependencies modeling, respectively. (c) Illustration of the discriminative query propagation, which learns more discriminative queries to represent high-level object information.
@@ 3.2 Spatial-Semantic Feature Learning
Figure 2(b) shows the spatial-semantic network, which consists of a semantic embedding module that incorporates semantic information from a trained ViT model into multi-scale features and a spatial dependency modeling module that learns spatial dependencies based on these incorporated features.
Semantic embedding. As the VOS task is designed for generic objects and no class labels are given, it is challenging to learn semantic representations directly from the VOS dataset during training. Fortunately, the CLS token in a trained ViT model aggregates semantic information from the entire image and provides a rich, global representation of the image content. Thus, we incorporate the CLS token with the multi-scale features generated from CNN networks to obtain detailed semantic features at different scales. To be specific, we use a cross-attention operation to perform the corporation of these two features by taking the semantic feature (CLS token $\mathcal{F}_{cls}^{i}$) concatenated with global token representation ( $\mathcal{F}_{g}^{i}$) as the Key and Value, and the spatial-temporal features ($\mathcal{F}_{spsem}^{i-1}$) from the $(i-1)$-th block as the Query to produce the semantic-aware features $\mathcal{F}_{sem}^{i}$ as:
@@
Spatial dependency modeling. We note that it is important to learn spatial dependencies between semantic features for the VOS model to understand the relationships between different object parts, which helps handle objects with complex structures or separate parts. Given the features of the input image patches generated by the ViT backbone model $\mathcal{F}_{vit}^{i}$, we adopt the Multi-scale Deformable Cross-Attention operation [5] to capture the spatial information ${\mathcal{F}}_{spa}^{i}$:
@@
where $\Delta p$ indicates the deformable offsets generated based on the query features $\mathcal{F}_{vit}^{i}$, $\textit{f}_{gs}$ performs grid sampling on $\mathcal{F}_{sem}^{i}$ using the offsets $\Delta p$, $\textit{f}_{\alpha}$ denotes the attention weights generated based on the query features $\mathcal{F}_{vit}^{i}$ and the grid sampled features, and $\sum_{j}$ is the summation over all sampled points.
To enable an effective long-range propagation through the cascade blocks, we use the residual connection [21] followed by a convolution-feed-forward to generate the final spatial-semantic feature of the current block:

### v2

@@ Abstract
Tracking and segmenting multiple similar objects with distinct or complex parts in long-term videos is particularly challenging due to the ambiguity in identifying target components and the confusion caused by occlusion, background clutter, and changes in appearance or environment over time. In this paper, we propose a robust video object segmentation framework that learns spatial-semantic features and discriminative object queries to address the above issues. Specifically, we construct a spatial-semantic block comprising a semantic embedding component and a spatial dependency modeling part for associating global semantic features and local spatial features, providing a comprehensive target representation. In addition, we develop a masked cross-attention module to generate object queries that focus on the most discriminative parts of target objects during query propagation, alleviating noise accumulation to ensure effective long-term query propagation. Extensive experimental results show that the proposed method achieves state-of-the-art performance on benchmark data sets, including the DAVIS2017 test (87.8%), YoutubeVOS 2019 (88.1%), MOSE val (74.0%), and LVOS test (73.0%), and demonstrate the effectiveness and generalization capacity of our model. The source code and trained models are released at https://github.com/yahooo-m/S3.
@@ 1 Introduction
Video Object Segmentation (VOS) aims to track and segment target objects specified in the initial frame with mask annotations in a video sequence (Xu et al., 2018; Pont-Tuset et al., 2017; Voigtlaender et al., 2019; Yang et al., 2021). This practice holds significant promise in various applications, particularly as the prevalence of video content increases in domains such as autonomous driving, augmented reality, and interactive video editing (Park et al., 2022; Bao et al., 2018). The main challenges faced by VOS include drastic target appearance change, occlusion, and identity confusion caused by similar objects and background clutter, which becomes more difficult in long-term videos.
@@
Existing VOS methods (Oh et al., 2019; Cheng et al., 2021a) typically perform video object segmentation by comparing the test frame with past frames. Specifically, they first use an association model to generate the correlated features of the test sample and target templates. Then, the target masks are predicted based on the correlated features. The pixel-wise correlated features (Wu et al., 2023) contribute to accurate mask predictions. To account for targets that appear differently over time, some methods (Yang et al., 2021; Cheng & Schwing, 2022) utilize a memory module to store these changing appearances. In addition, several recent approaches (Wang et al., 2023; Cheng et al., 2023a) introduce object queries to help distinguish different target objects to alleviate identity confusion.
@@
Although significant advances have been made in video object segmentation, existing methods do not perform well in challenging scenarios. First, when dealing with target objects with multiple complex or separate parts caused by occlusion, background clutter, and shape complexity, existing methods often generate incomplete prediction masks. This is because the pixel-wise correlation adopted by existing methods mainly focuses on detailed spatial information at the pixel level and neglects the semantic information at the instance level. Second, although object query improves the ID association accuracy, it does not work well in sequences with dramatic target appearance changes. Existing methods (Wang et al., 2023; Cheng et al., 2023a) update target queries based on the entire predicted sample, which can introduce noise and errors and further lead to tracking failures when targets are missed or their identities are swapped.
@@
Figure 1: Performance comparison in terms of the $\mathcal{J}\&\mathcal{F}$ score on various datasets. The proposed method achieves favorable performance against state-of-the-art methods, including AOT_L (Yang et al., 2021), XMem (Cheng & Schwing, 2022), and Cutie (Cheng et al., 2023a) on all the datasets. All methods are trained using the YoutubeVOS and DAVIS datasets.
@@
To address the aforementioned issues, we propose a VOS algorithm, called S3, to learn spatial-semantic features and generate discriminative object queries for robust VOS. Specifically, we design a spatial semantic block with a semantic embedding module and a spatial dependencies modeling part. This block efficiently utilizes the semantic information and local details from pre-trained ViTs for VOS, avoiding the need to train all ViT backbone parameters. In addition, we propose a discriminative query that emphasizes representative target parts, enabling more reliable target representation and query updates. This approach mitigates noise accumulation during query propagation and enhances robustness in long-term videos. We evaluate the proposed method extensively on five benchmarks, including DAVIS 2017 (Pont-Tuset et al., 2017), Youtube VOS2018 (Xu et al., 2018), Youtube VOS2019, LVOS (Hong et al., 2023), and MOSE (Ding et al., 2023). Extensive experimental results show that our approach achieves state-of-the-art performance on all these datasets including the DAVIS2017 test (89.1%), YoutubeVOS 2019 (88.5%), MOSE (75.1%), and LVOS test (73.0%) datasets. In addition, the results achieved by only using the DAVIS and Youtube datasets for training show that our method generalizes well to different scenarios, as shown in Figure 1.
@@
  • We propose a spatial-semantic network block to incorporate semantic information with spatial information for Video Object Segmentation. The spatial-semantic block integrates global semantic information from the CLS token of a pre-trained ViT backbone into the base features of the input sample and then models spatial dependencies using a spatial dependency modeling module, enhancing robust VOS performance when handling target objects with complex or separate parts.
  • We develop a discriminative query mechanism to capture the most representative region of the target for better target representation learning and updating. This significantly improves VOS datasets, especially in long-term VOS datasets.
  • We demonstrate that the proposed method achieves state-of-the-art performance on five diverse datasets and evaluate the contribution of each proposed component with comprehensive ablation studies.
@@ 2 Related Work
Video Object Segmentation. Video Object Segmentation (VOS) aims to separate objects from the background and identify each target object across frames based on the target mask annotations given in the initial frame. Early VOS methods (Caelles et al., 2017; Xiao et al., 2018; Voigtlaender & Leibe, 2017; Hu et al., 2017; Luiten et al., 2018) use test-time learning to adapt pre-trained segmentation models to separate the specified targets online. The test-time learning of a deep model is time-consuming and significantly reduces the run speed. To avoid online learning, propagation-based methods (Duke et al., 2021; Cheng et al., 2018; Tsai et al., 2016) use offline shift attention to model temporal correlations between adjacent frames, propagating target masks from the previous frame to the current one. Despite promising performance, these methods suffer from error accumulation caused by inaccurate predictions. Matching-based methods (Chen et al., 2018; Hu et al., 2018; Bhat et al., 2020; Yang et al., 2020) enable fast and efficient inference by comparing the target template with the test image and predicting target masks based on feature matching. For example, OML (Chen et al., 2018) and VideoMatch (Hu et al., 2018) perform pixel matching between adjacent frames for mask prediction. Several methods (Oh et al., 2019; Yang et al., 2021; Yang & Yang, 2022; Cheng & Schwing, 2022) improve the matching-based VOS framework by introducing target memory to enrich target templates and developing advanced matching strategies to generate more comprehensive correlated features. For example, SimVOS (Wu et al., 2023) and JointFormer (Mao et al., 2021) use ViT blocks (Dosovitskiy et al., 2021) to model features and patch correspondence for object modeling jointly. For better target association across frames, some recent methods (Wang et al., 2023; Cheng et al., 2023a) introduce object queries to enrich target representation. Despite good performance on simple target objects, existing methods lack modeling of object-level semantic information, so they do not work well on dealing with objects with complex structures and long-term deformation. Unlike existing methods, the proposed approach performs pixel- and object-level feature learning through a novel object-aware backbone and a query-level correlation mechanism, enabling robust VOS performance, particularly in complex and long-term video scenarios.
@@
Query-Based VOS Frameworks. DETR (Carion et al., 2020) introduces the transformer network (Dosovitskiy et al., 2021) to object detection tasks by using queries to represent target regions. In the segmentation domain, some methods (Cheng et al., 2022; Cavagnero et al., 2024) introduce masked cross-attention modules to use queries to represent targets for segmentation tasks. Motivated by the above methods, ISVOS (Wang et al., 2023) uses additional queries generated by Mask2Former (Cheng et al., 2022) to represent instance information for target modeling, enhancing target representation in the matching-based VOS framework. To enable end-to-end training, Cutie (Cheng et al., 2023a) learns object-level features and enhances the interaction between object-level and pixel-level information, leading to improved segmentation accuracy and efficiency. However, query-based VOS methods suffer from query drift due to ineffective propagation. These methods model the entire target sample in the query, incorporating redundant information that diminishes the saliency of the target object. Instead, the proposed method generates discriminative queries for target representation and propagation, which preserves key distinguishing information of target objects and offers greater robustness to variations.
@@
Adapters for Dense Tasks. The ViT-Adapter (Chen et al., 2022) efficiently leverages pre-trained models by introducing a dense adapter for enhanced information exchange within ViT blocks, improving detection and segmentation performance. To enhance spatial hierarchical features, ViT-CoMer (Xia et al., 2024) introduces a multi-receptive field feature pyramid module to provide multi-scale spatial features. However, both methods require a complete fine-tuning step, which is inefficient, particularly in the video domain. Meanwhile, they focus on spatial information enhancement while lacking semantic feature interaction between ViT blocks and multi-scale features.
@@
To better adapt to VOS tasks, the proposed method improves adapter-based VOS frameworks in the following aspects. First, we introduce a spatial-semantic module to integrate the global semantic features from ViT to multi-scale features. This module does not need extra modules to transmit information to ViT blocks, which is more efficient. Second, the proposed spatial-semantic integration block associates semantic priors with dense local features. Third, the proposed method freezes the ViT blocks during training, eliminating the need for full fine-tuning.
@@ 3 Proposed Algorithm
Our method aims to learn comprehensive target features that contain semantic, spatial, and discriminative information for video object segmentation. This helps us cope with complex target appearance variations and ID confusion between target objects with similar appearances in long-term videos. To this end, we propose a spatial-semantic feature learning network that first embeds semantic information from a trained ViT network with multi-scale features from a trainable CNN network. Then, we model spatial dependencies upon the fused feature, enabling spatial-semantic features for VOS. In addition, to ensure that the target query retains discriminative information during long-term appearance changes, we develop a discriminative query propagation module to capture and update local representations of the target object.
@@ 3.1 Overall Framework
Figure 2(a) shows the overall framework of the proposed method, consisting of two core modules: a spatial-semantic feature generation module and a pixel-and-query dual-level target association module. Given a test frame $\mathbf{X}_{t}\in\mathbb{R}^{3\times H\times W}$, the proposed method predicts the mask of all specified objects within it based on the initial frame $\mathbf{X}_{init}\in\mathbb{R}^{3\times H\times W}$ and its corresponding annotation mask $\mathbf{M}_{init}\in\mathbb{R}^{H\times W}$. First, the feature generation module takes the test frame $\mathbf{X}_{t}$ as input and generates spatial-semantic features. Then, the target association module associates the spatial-semantic features with the reference samples pixel-wise and query-wise to better represent the correlated features. Finally, a decoder predicts the mask based on these correlated features. In addition, we develop a dual-level memory module consisting of target features and target queries to better represent video targets that vary over frames. Both kinds of memory are updated every few frames using the online predicted masks, enabling adaptive target representations for effective target association.
Figure 2: Architecture of the proposed method. (a) shows the overall framework of the proposed method comprising a feature generation part, a target association component, and a decoder-based prediction part. The feature generation part contains multiple spatial-semantic blocks, as illustrated in (b), which learn spatial-semantic features by integrating semantic priors and spatial details. (c) illustrates the discriminative query propagation process, which learns to generate discriminative queries that represent high-level object information.
@@ 3.2 Spatial-Semantic Feature Learning
Figure 2(b) shows the spatial-semantic network, which consists of a semantic embedding module that incorporates semantic information from a pre-trained ViT model into multi-scale features and a spatial dependency modeling module that learns spatial dependencies based on semantic features.
Semantic embedding. As the VOS task is designed for generic objects and no class labels are given, it is challenging to learn semantic representations directly from the VOS dataset during training. Fortunately, the CLS token in a trained ViT model aggregates semantic information from the entire image and provides a rich, global representation of the image content. Thus, we incorporate the semantic features generated from the CLS token with the multi-scale features generated from CNN networks to obtain detailed semantic features at different scales. To be specific, we first concatenate the CLS token $\mathcal{F}_{cls}^{i}$ with a global token representation $\mathcal{F}_{g}$, which is generated by averagely pooling the features from a ViT block, to generate global semantic features denoted as ($\mathcal{F}_{cls}^{i}$, $\mathcal{F}_{g}$). Then, we use a cross-attention operation to generate the semantic-aware features $\mathcal{F}_{sem}^{i}$ by taking ($\mathcal{F}_{cls}^{i}$, $\mathcal{F}_{g}$) as the Key and Value, and the spatial-temporal features ($\mathcal{F}_{spsem}^{i-1}$) from the $(i-1)$-th spatial-semantic block as the Query, which is formulated as:
@@
Spatial dependency modeling. We note that learning spatial dependencies between semantic features is important for the VOS model to understand the relationships between different object parts, which helps handle objects with complex structures or separate parts. Given the features of the input image patches generated by a ViT backbone model $\mathcal{F}_{vit}^{i}$, we adopt the Multi-scale Deformable Cross-Attention operation (Carion et al., 2020) to capture the spatial information ${\mathcal{F}}_{spa}^{i}$ as:
@@
where $\Delta p$ indicates the deformable offsets generated based on the query features $\mathcal{F}_{vit}^{i}$, $\textit{f}_{gs}$ performs grid sampling on $\mathcal{F}_{sem}^{i}$ using the offsets $\Delta p$, $\textit{f}_{\alpha}$ denotes the attention weights generated based on the query features $\mathcal{F}_{vit}^{i}$ and the grid sampled features, and $\sum_{j}$ is the summation over all sampled points. To enable an effective long-range propagation through the cascade blocks, we use the residual connection (He et al., 2016) followed by a convolution-feed-forward to generate the final spatial-semantic feature of the current block:
@@
Figure 3: Visualized feature maps from different backbones and stages. It shows that the proposed spatial-semantic model generates effective features for target representation.
Figure 3(a) shows how the feature changes during the processing of the proposed model by visualizing the feature maps in different processing steps. It shows that the semantic embedding module enhances the semantic information (boundaries) of the objects in the feature maps. In contrast, after spatial dependence modeling, the feature maps pay more attention to the details of the objects. Figure 3(b) compares the features generated by different networks. Compared to ViT and ResNet models, the proposed Spatial-Semantic model exhibits better target representation capabilities with significantly higher activations in target object areas.